OpenStudy (anonymous):

–3(–4y + 3) + 7y Combine like terms

OpenStudy (anonymous):

is the answer 19y^2 - 3 or 19y -9 ?

OpenStudy (debbieg):

Which do you think? :)

OpenStudy (anonymous):

those are both the answers I came up with... i just solved the problem two different ways hahaha I'm not sure which way is the right way to solve it

OpenStudy (debbieg):

OK, well.... one of the ways you did it is right and one is wrong. :) Here is the right way: –3(–4y + 3) + 7y distribute = 12y - 9 + 7y put like terms together = 12y + 7y -9 add like terms = 19y - 9 You can't get a y^2 term from there, because there is NOWHERE that you are multiplying that y by another factor of y. And I have no idea where the -3 would come from either.... so whatever you did to get THAT answer is wrong. If you want to show me what you did, I can walk you through where your error is.

OpenStudy (anonymous):

I distributed the -3 at the beginning

OpenStudy (debbieg):

ok, that's correct - you should get 12y - 9 + 7y and then?

OpenStudy (anonymous):

Ohhhhhh I see what I did! I forgot to distribute the -3 to the positive 3 lol

OpenStudy (debbieg):

OK, that still doesn't explain how you had a -3 instead of 3 (I guess you distributed the "-" but not the 3? lol) or how you got the y^2... you can't get y^2 from y, unless you multiply by another y. :) But as long as you see what you did wrong and won't do it next time, that's all that matters!! :)

OpenStudy (debbieg):

Oh, I think you added 12y + 7y and made it 19y^2? but you see why that doesn't work now, right? :) When you add like terms, you JUST add together the coefficients!

OpenStudy (anonymous):

Haha yes thats exactly what I did.. thank you :)

OpenStudy (debbieg):

you're welcome. :)